OpenStudy (anonymous):

the light bulbs you use are sale. But you only get the lower price if you buy three packages of bulbs you want to know how much you save. Write a formula that will help you determine how much you'll save by buying 3 packages on sale let's let s=sales price per package r=regular price per package c=cost savings

OpenStudy (anonymous):

what you want to do is find out the cost savings. So what you have to do is get the regular prices for buying 3 bulbs, then get the sales prices of 3 bulbs. The cost savings lies in the difference. ;) Good Luck

OpenStudy (anonymous):

i have no idea what to do :/

OpenStudy (anonymous):

What he said is that you need to find out the normal price for buying 3 bulbs. 3r r=regular price per package Then he said what the SALES price for buying 3 bulbs is. If s=sales price per package, then what is it now?

OpenStudy (anonymous):

then what is what?

OpenStudy (anonymous):

what is the equation for sales price, Hint:it's a LOT like the equation for regular price per package, just change a variable.

OpenStudy (anonymous):

umm 3?

OpenStudy (anonymous):

what is the variable for the sale price per package?

OpenStudy (anonymous):

umm 6?

OpenStudy (anonymous):

do you know what a variable is?

OpenStudy (anonymous):

no lol

OpenStudy (anonymous):

ah ok, a variable is a letter that can be any number, so we represent it with a letter. Such as r, can be a variable in 3r. r can be ANY number in the whole wide world and you can stick it into that slot. We choose it for formulas and such. r is the variable for regular price per package. So what is s?

OpenStudy (anonymous):

how would i figure out is

OpenStudy (anonymous):

s*

OpenStudy (anonymous):

You don't, what I'm asking is what does the variable s stand for in this problem?

OpenStudy (anonymous):

sales price per package

OpenStudy (anonymous):

great job! Ok, what does the variable c stand for?

OpenStudy (anonymous):

cost savings

OpenStudy (anonymous):

alrighty, so now do you understand a variable? Do you need any more help on that?

OpenStudy (anonymous):

i just needa kno how to find the answer cause im takin a test

OpenStudy (anonymous):

ok, 3r is the equation for the regular price of 3 bulbs. since we know that, what is the sales price of 3 bulbs equation?

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

no, it is an equation, not a number.

OpenStudy (anonymous):

a formula, what is the formula for the sales price of 3 bulbs.

OpenStudy (anonymous):

wdym formula

OpenStudy (anonymous):

A mathematical relationship or rule expressed in symbols.

OpenStudy (anonymous):

For example, the Pythagorean Theorem is a formula: \[A^2+B^2=C^2\]

OpenStudy (anonymous):

ok so am i trying the formula?

OpenStudy (anonymous):

right, we're building to that. I've given you the formula for finding the regular price per package for 3 bulbs. It is \[3r\]

OpenStudy (anonymous):

3r s/c

OpenStudy (anonymous):

we still have to get the formula for the sales price of 3 bulbs.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

What does s stand for?

OpenStudy (anonymous):

sales price per package

OpenStudy (anonymous):

alright! So now that you know that, we're trying to make a formula for the sales price per package for 3 bulb packages.

OpenStudy (anonymous):

so that would be 3s, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okey doke. Ok,here's a bit of common sense. When you get a sale on something, you get some money taken off, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, so the regular price is more, and we're trying to figure out how much we save by getting the bulbs on sale. so what would the order of both formulas be? The formulas are 3r and 3s

OpenStudy (anonymous):

you're subtracting one from the other

OpenStudy (anonymous):

subtrcting what from the others

OpenStudy (anonymous):

you're subtracting either 3r - 3s or 3s - 3r Remember what the variables stand for and what we're trying to get in the end!

OpenStudy (anonymous):

so would it just be 0

OpenStudy (anonymous):

no, remember it is a formula. We're trying to figure out which way the formula is going to subtract.

OpenStudy (anonymous):

is it 3r-3s=c OR 3s-3r=c

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

great job!! You did really well :)

OpenStudy (anonymous):

thanks! (:

OpenStudy (anonymous):

your welcome, and do you think that you'll understand variables a bit more from now on?

OpenStudy (anonymous):

ooh idk! ill try

OpenStudy (anonymous):

if you need any help, you can ask me or ask on here.

OpenStudy (anonymous):

the website, I mean, not this question.

OpenStudy (anonymous):

okay! ;] but was the answer?

OpenStudy (anonymous):

3r-3s=c

OpenStudy (anonymous):

thx